OpenStudy (anonymous):

In a random sample of 50 undergraduate students at a college, it was found that 44 students regularly access social networking websites from their college library. What is the margin of error for the true proportion of all undergraduates who access social networking sites from their college library?

OpenStudy (amistre64):

margins of error tend to have a confidence interval associated with them

OpenStudy (amistre64):

you will need to know the sample propotion, and its complement; as well as the sample size

OpenStudy (aggie12341):

that many students don't know the process of expiration dates

OpenStudy (amistre64):

if we assume a test statistic of 1 .. then the margin of error can be calculated as:\[\sqrt{\frac{x(n-x)}{n^3}}\]

OpenStudy (anonymous):

I have not the slightest a' clue as to how to approach this, I've never even seen the equation you've written down. If you could somehow manage to explain to me how I can solve this that would be great. @amistre64 @aggie12341

OpenStudy (anonymous):

I've tried plugging in 50 and 44 into the equation. x=50, n=44. Then I flipped them; n=50, x=44, and I still didn't get my answer that was listed.

OpenStudy (amistre64):

from the equation to find a z score: \[z=\frac{x-\bar x}{s/\sqrt n}\] \[z\frac{s}{\sqrt n} =x-\bar x\] \[\bar x+ z\frac{s}{\sqrt n} =x\] this is the approach to a confidence interval with a margin of error of that z part

OpenStudy (amistre64):

it can be shown that \(\large s = \sqrt{\frac xn~\frac{n-x}{n}}\)

OpenStudy (amistre64):

what are the options you have?

OpenStudy (anonymous):

1.)0.046 2.)0.038 3.)0.014 4.)0.070 5.)0.092

OpenStudy (amistre64):

i get one of them ... what results did you get when you plugged in for: x = 44, n = 50

OpenStudy (anonymous):

-0.00049805605361561856057535055557723213698687730396915432704...

OpenStudy (amistre64):

http://www.wolframalpha.com/input/?i=sqrt%2844%2850-44%29%2F50%5E3%29 you must be entering it in wrong .. or hitting the wrong buttons

OpenStudy (anonymous):

I've tried using wolframalpha, it won't allow all the numbers to go under the square root.

OpenStudy (amistre64):

you have to use your paranthesis correctly sqrt() sqrt(44(50-44)/50^3) works fine

OpenStudy (anonymous):

It was using sqrt 50 * the rest of the equation.

OpenStudy (anonymous):

Oh, that might have been the mistake, thank you.

OpenStudy (amistre64):

good luck :)